### v3

# Ensemble Kalman Filtering Meets Gaussian Process SSM for Non-Mean-Field and Online Inference

Abstract
Gaussian process state-space models (GPSSMs) are a versatile and principled family of nonlinear dynamical system models. However, existing variational learning and inference methods for GPSSMs often necessitate optimizing a substantial number of variational parameters, leading to inadequate performance and efficiency. To overcome this issue, we propose incorporating the ensemble Kalman filter (EnKF), a well-established model-based filtering technique, into the variational inference framework to approximate the posterior distribution of latent states. This utilization of EnKF can effectively exploit the dependencies between latent states and GP dynamics, while eliminating the need for parameterizing the variational distribution, thereby significantly reducing the number of variational parameters. Moreover, we show that our proposed algorithm allows straightforward evaluation of an approximated evidence lower bound (ELBO) in variational inference via simply summating multiple terms with readily available closed-form solutions. Leveraging automatic differentiation tools, we hence can maximize the ELBO and train the GPSSM efficiently. We also extend the proposed algorithm to accommodate an online setting and provide detailed algorithmic analyses and insights. Extensive evaluation on diverse real and synthetic datasets demonstrates the superiority of our EnKF-aided variational inference algorithms in terms of learning and inference performance compared to existing methods.
Index Terms:
Gaussian process, state-space model, ensemble Kalman filter, online learning, variational inference.
I Introduction
State-space models (SSMs) describe the underlying dynamics of latent states through a transition function and an emission function [1]. As a versatile tool for modeling dynamical systems, SSM finds successful applications in diverse fields, including control engineering, signal processing, computer science, and economics [2, 3, 4]. In SSM, a key task is to infer unobserved latent states from a sequence of noisy measurements. Established techniques, such as the Kalman filter (KF), extended Kalman filter (EKF), ensemble Kalman filter (EnKF), and particle filter (PF), have been widely employed over the decades for latent state inference [2]. However, these classic state inference methods heavily rely on precise knowledge of the underlying system dynamics [5], posing tremendous challenges in complex and uncertain scenarios like model-based reinforcement learning [6] and disease epidemic propagation [7]. As a novel alternative, the underlying complex dynamics can be learned from noisy observations, leading to the emergence of data-driven SSMs. One class of prominent data-driven SSMs is the Gaussian process state-space model (GPSSM) [8], which utilizes Gaussian processes (GPs) [9] as the core learning modules to capture the complex underlying system dynamics.
Gaussian processes, serving as the most prominent Bayesian nonparametric model [10], provide the flexibility to model general nonlinear system dynamics without enforcing an explicit parametric structure. With the inherent regularization imposed by the GP prior, GPSSMs are able to mitigate model overfitting and generalization issues [11], rendering them to be more effective in scenarios with limited data samples [12]. Moreover, owing to its Bayesian nature, GPSSMs maintain good interpretability and explicit uncertainty calibration for analyzing system dynamics. These superior properties have led to the extensive usage of GPSSMs in various live applications, such as human pose and motion learning [13], robotics and control learning [14], reinforcement learning [15], target tracking and navigation [16], and magnetic-field sensing [17].
Despite the popularity of GPSSMs, simultaneously learning the model and estimating the latent states in GPSSMs remains highly challenging, primarily due to the following two factors. First, the quality of the state estimates affects the model learning and vice versa, leading to heightened computational and statistical complexities. Second, the nonlinearity in GPSSMs prohibits tractable learning and inference process [10]. Hence, the main task in GPSSM is to accurately approximate the joint posterior over the latent states and the system dynamics represented by GPs. Multiple approaches have emerged towards achieving this objective in the last decade.
The seminal work utilizing particle Markov chain Monte Carlo (PMCMC) was proposed in [8]. Subsequently, more advanced methods [18, 19, 20, 21] leveraged the reduced-rank GP approximation introduced in [22] to alleviate the substantial computational demands of GPs in [8]. However, the computational burden of the involved PMCMC remains unaffordable, particularly when dealing with high-dimensional latent states. Consequently, there has been a paradigm shift towards variational inference methods [23, 24, 25, 26, 27, 28, 29, 30, 31], which adopted the classic sparse GP approximations with inducing points [32].
Generally, the variational inference approaches can be classified into two main categories: the mean-field (MF) class [23, 24, 25, 26] versus the non-mean-field (NMF) class [27, 28, 29, 30, 31], depending on whether the statistical independence assumption has to be applied to the variational distribution of the latent states and GP dynamics. The first variational algorithm, incorporating the MF assumption and integrating PF for latent state inference, was introduced in [23]. Subsequent MF endeavors [24, 25, 26] aimed at reducing computational and model complexities in [23]. Yet, the main idea of the MF class is to neglect the inherent dependencies between the latent states and the GP dynamics, potentially impairing the overall learning accuracy and yielding overconfident state estimates [33].
Recent methods [27, 28, 29, 30, 31], therefore, concentrate on the more complicated NMF approximation to address the inaccuracies arising from the MF assumption, yet they continue to struggle with computational challenges. Specifically, in [27], the posterior distribution of underlying states conditioned on the GP dynamics was simply approximated by the subjectively selected prior distribution, while in [28] and [29], the posterior was approximated by some parametric Markov-structured Gaussian variational distributions. However, optimizing the extensive variational parameters in the variational distributions is computationally demanding and inefficient. In contrast to the heavy parameterization as conducted in [27, 28, 29], the works in [30] and [31] opted for directly characterizing the posterior distribution through the model evidence. Nonetheless, the Laplace approximation, employed in [30], can yield an oversimplified unimodal posterior distribution over the latent states [10], while the stochastic gradient Hamiltonian Monte Carlo method [34], employed in [31], may exhibit very slow convergence ($50,000$ iterations as reported). Furthermore, all these existing variational methods are incompatible with online applications, posing a notable limitation in their applicability.
To tackle the computational challenges mentioned above, our proposed methods incorporate the model-based filtering technique, EnKF, to approximate the joint posterior distribution. By leveraging the EnKF, our method can fully exploit the dependencies between the latent states and the GP dynamics, while reducing the parameterization of associated variational distributions, thereby alleviating model and computational complexities and accelerating algorithm convergence. This streamlined parameterization also helps the easy accommodation of our method in an online setting. The main contributions of this paper are summarized as follows.
  • We corroborate the significance of incorporating well-established model-based filtering techniques to enhance the efficiency of GPSSM variational algorithms. In contrast to existing methods utilizing numerous variational parameters to parameterize the variational distribution over latent states, our proposed novel algorithm harnesses the EnKF to accurately capture the dependencies between the latent states and the GP dynamics, and eliminate the need to parameterize the variational distribution, significantly reducing the number of variational parameters and accelerating the algorithm convergence.
  • By leveraging EnKF, we also demonstrate a simplified evidence lower bound (ELBO) for variational inference via simply summating multiple terms with readily available closed-form solutions. Leveraging the differentiable nature of the classic EnKF alongside the off-shelf automatic differentiation tools, we can optimize the ELBO and train the GPSSM efficiently.
  • Without explicitly parameterizing the variational distribution over latent states, the proposed EnKF-aided algorithm gets readily extended to accommodate online applications. This online method relies on a more principled objective by ensuring observation fitting accuracy while incorporating model regularizations to mitigate model overfitting, rendering it superior to state-of-the-art online algorithms in learning and inference efficiency.
  • We provide detailed analysis and fresh insights for the proposed algorithms and conduct extensive experiments on various real and synthetic datasets to evaluate the performance of these algorithms from various aspects. The results demonstrate that the proposed EnKF-aided variational learning methods consistently outperform the existing state-of-the-art methods, especially in terms of learning and inference.
The remainder of this paper is organized as follows. Some preliminaries related to GPSSMs are provided in Section II. The computational challenges of the existing methods are summarized in Section III. Section IV elaborates the proposed EnKF-aided variational learning algorithm. Section V extends the proposed algorithm to accommodate online applications. The experimental results are presented in Section VI, and Section VII concludes this paper. More supportive results, proofs, etc. are relegated to the Appendix and supplementary materials.
II Preliminaries
In Section II-A, we briefly review the Gaussian process regression. Section II-B is dedicated to introducing Gaussian process state-space models.
II-A Gaussian Processes (GPs)
A GP defines a collection of random variables indexed by $\bm{x}\in\mathcal{X}$, such that any finite subset of these variables follows a joint Gaussian distribution [9]. Mathematically, a real scalar-valued GP $f(\bm{x})$ can be represented as
$f(\bm{x})\sim\mathcal{GP}\left(\mu(\bm{x}),\ k(\bm{x},\bm{x}^{\prime});\ \bm{\theta}_{gp}\right),$ (1)
where $\mu({\bm{x}})$ is a mean function typically set to zero in practice, and $k(\bm{x},\bm{x}^{\prime})$ is the kernel function that provides insights about the nature of the underlying function [9]; and $\bm{\theta}_{gp}$ is a set of hyperparameters that needs to be tuned for model selection. By placing a GP prior over the function $f(\cdot):{\cal X}\mapsto\mathbb{R}$ in a general regression model,
$y=f(\bm{x})+{e},\quad{e}\sim{\cal N}(0,\sigma_{e}^{2}),\quad y\in\mathbb{R},$ (2)
we get the salient Gaussian process regression (GPR) model.
Given an observed dataset, $\mathcal{D}\!\triangleq\!\{\bm{x}_{i},y_{i}\}_{i=1}^{n}\!\triangleq\!\{\bm{X},\bm{y}\}$ consisting of $n$ input-output pairs, the posterior distribution of the mapping function, $p(f(\bm{x}_{*})|\bm{x}_{*},{\cal D})$, at any test input $\bm{x}_{*}\!\in\!{\cal X}$, is Gaussian [9], fully characterized by the posterior mean $\xi$ and the posterior variance $\Xi$. Concretely,
$\!\xi(\bm{x}_{*})\!=\!\bm{K}_{\bm{x}_{*},\bm{X}}\left(\bm{K}_{\bm{X},\bm{X}}+\sigma_{e}^{2}\bm{I}_{n}\right)^{\!-\!1}{\bm{y}},$ (3a)
$\!\Xi(\bm{x}_{*})\!=\!k(\bm{x}_{*},\bm{x}_{*})\!-\!\bm{K}_{\bm{x}_{*},\bm{X}}\left(\bm{K}_{\bm{X},\bm{X}}\!+\!\sigma_{e}^{2}\bm{I}_{n}\right)^{\!-\!1}\bm{K}_{\bm{x}_{*},\bm{X}}^{\top},$ (3b)
where $\bm{K}_{\bm{X},\bm{X}}$ denotes the covariance matrix evaluated on the training input $\bm{X}$, and each entry is $[\bm{K}_{\bm{X},\bm{X}}]_{i,j}=k({\bm{x}}_{i},\bm{x}_{j})$; $\bm{K}_{\bm{x}_{*},\bm{X}}$ denotes the cross covariance matrix between the test input $\bm{x}_{*}$ and the training input $\bm{X}$; the zero-mean GP prior is assumed here and will be used in the rest of this paper if there is no further specification. Note that the posterior distribution $p(f(\bm{x}_{*})|\bm{x}_{*},{\cal D})$ gives not only a point estimate, i.e., the posterior mean, but also an uncertainty region of such estimate quantified by the posterior variance. It is also noteworthy that here we denote the variables in the GPR using mathematical mode italics, such as $\bm{x}_{i}$ and $y_{i}$; these variables should not be confused with the latent state ${\mathbf{x}}_{t}$ and observation ${\mathbf{y}}_{t}$ in SSM (cf. Eq. (II-B)).
II-B Gaussian Process State-Space Models (GPSSMs)
A generic SSM describes the probabilistic dependence between latent state ${\mathbf{x}}_{t}\in\mathbb{R}^{d_{x}}$ and observation ${\mathbf{y}}_{t}\in\mathbb{R}^{d_{y}}$. Mathematically, it can be expressed by the following equations:
[…]
where the latent states form a Markov chain. That is, for any time instance $t\!\in\!\mathbb{N}$, the next state ${\mathbf{x}}_{t+1}$ is generated by conditioning on only ${\mathbf{x}}_{t}$ and the transition function $f(\cdot)\!:\!\mathbb{R}^{d_{x}}\!\mapsto\!\mathbb{R}^{d_{x}}$. The emission is assumed to be linear with a known coefficient matrix, $\bm{C}\in\mathbb{R}^{d_{y}\times d_{x}}$, hence mitigating the system non-identifiability111 Nonlinear emissions can be addressed by augmenting the latent state to a higher dimension. This augmentation helps mitigate/eliminate the significant non-identifiability issues commonly encountered in GPSSMs [35, 24]. The initial state distribution $p({\mathbf{x}}_{0})$ is also known and assumed to follow a Gaussian distribution. Both the states and observations are corrupted by zero-mean Gaussian noise with covariance matrices $\mathbf{Q}$ and $\mathbf{R}$, respectively.
The GPSSM incorporates a GP prior to model the transition function $f(\cdot)$ in Eq. (II-B). Specifically, Fig. 1 presents the graphical model of GPSSM, while the following equations express its mathematical representation:
$f(\cdot)\sim\mathcal{G}\mathcal{P}\left(\mu(\cdot),k(\cdot,\cdot);{{\bm{\theta}}}_{gp}\right)$ (5a)
$\mathbf{x}_{0}\sim p\left(\mathbf{x}_{0}\right)$ (5b)
${{\mathbf{f}}}_{t}=f\left(\mathbf{x}_{t-1}\right)$ (5c)
$\mathbf{x}_{t}\mid{{\mathbf{f}}}_{t}\sim\mathcal{N}\left({\mathbf{x}}_{t}\mid{{\mathbf{f}}}_{t},\mathbf{Q}\right)$ (5d)
$\mathbf{y}_{t}\mid\mathbf{x}_{t}\sim{\cal N}\left(\mathbf{y}_{t}\mid\bm{C}\mathbf{x}_{t},\mathbf{R}\right)$ (5e)
where $\mathbf{f}_{t}$ represents the GP transition function value evaluated at the previous state $\mathbf{x}_{t-1}$. Note that when dealing with a multidimensional state space ($d_{x}\!>\!1$), the transition function $f(\cdot):\mathbb{R}^{d_{x}}\!\mapsto\!\mathbb{R}^{d_{x}}$ is represented using a multi-output GP. In this context, the $d_{x}$ functions are typically modeled with $d_{x}$ mutually independent GPs [36]. Furthermore, it is noteworthy that the GPSSM illustrated in Fig. 1 can be extended to accommodate control systems incorporating a deterministic control input $\bm{c}_{t}\in\mathbb{R}^{d_{c}}$ by augmenting the latent state with $[{\mathbf{x}}_{t},\bm{c}_{t}]\in\mathbb{R}^{d_{x}+d_{c}}$. For the sake of brevity, however, we omit explicit reference to $\bm{c}_{t}$ in our notation throughout this paper.
${{\mathbf{x}}}_{0}$ ${{\mathbf{x}}}_{1}$cdots $\!\!{{\mathbf{x}}}_{t-1}\!\!$ ${{\mathbf{x}}}_{t}$cdots $\cdots$ ${{\mathbf{f}}}_{1}$cdots $\!\!{{\mathbf{f}}}_{t-1}\!\!$ ${{\mathbf{f}}}_{t}$cdots ${{\mathbf{y}}}_{1}$cdots $\!\!{{\mathbf{y}}}_{t-1}\!\!$ ${{\mathbf{y}}}_{t}$cdots
Figure 1: Graphical model of GPSSM. The white circles represent the latent variables, while the gray circles represent the observable variables. The thick horizontal bar represents a set of fully connected nodes, i.e., the GP.
Given the aforementioned model descriptions, the joint density function of the GPSSM can be expressed as:
$p({\vec{\mathbf{f}}},{{\vec{\mathbf{x}}}},{\vec{\mathbf{y}}}|\bm{\theta})=p(\mathbf{x}_{0})p({\mathbf{f}}_{1:T})\prod_{t=1}^{T}p(\mathbf{y}_{t}|\mathbf{x}_{t})p(\mathbf{x}_{t}|{{\mathbf{f}}}_{t}),$ (6)
where $p({\mathbf{f}}_{1:T})\!\!=\!\!p(f({\mathbf{x}}_{0:T-1}))\!\!=\!\!\prod_{t=1}^{T}p({\mathbf{f}}_{t}|{\mathbf{f}}_{1:t-1},{\mathbf{x}}_{0:t-1})$ corresponds to a finite dimensional ($T$-dimensional) GP distribution [24], and we define the short-hand notations ${\vec{\mathbf{y}}}\triangleq{\mathbf{y}}_{1:T}=\{{\mathbf{y}}_{t}\}_{t=1}^{T}$, ${\vec{\mathbf{f}}}\triangleq{\mathbf{f}}_{1:T}=\{{\mathbf{f}}_{t}\}_{t=1}^{T}$, and ${\vec{\mathbf{x}}}\triangleq{\mathbf{x}}_{0:T}=\{{\mathbf{x}}_{t}\}_{t=0}^{T}$. The model parameters $\bm{\theta}$ includes the noise covariance and GP hyper-parameters, i.e., $\bm{\theta}\!\!=\!\![\mathbf{Q},\mathbf{R},\bm{\theta}_{gp}]$. The challenging task in GPSSM is to learn ${{\bm{\theta}}}$, and simultaneously infer the latent states of interest, which involves the marginal distribution $p({\vec{\mathbf{y}}}|{{\bm{\theta}}})$. However, due to the nonlinearity of GP, a closed-form solution for $p({\vec{\mathbf{y}}}|{{\bm{\theta}}})$ is unavailable. Hence, it becomes necessary to utilize approximation methods.
III Problem Statement
To overcome the intractability of the marginal distribution $p({\vec{\mathbf{y}}}|\bm{\theta})$, variational GPSSMs involve constructing a model evidence lower bound (ELBO) to the logarithm of the marginal likelihood. Specifically, the ELBO, denoted by $\mathcal{L}$, is constructed such that the difference between $\log p({\vec{\mathbf{y}}}|{{\bm{\theta}}})$ and $\mathcal{L}$ is equal to the Kullback-Leibler (KL) divergence between the variational approximation, $q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})$, and the true posterior, $p({\vec{\mathbf{x}}},{\vec{\mathbf{f}}}|{\vec{\mathbf{y}}})$, i.e.,
$\log p({\vec{\mathbf{y}}}|{{\bm{\theta}}})-\underbrace{\mathbb{E}_{q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})}\!\left[\log\frac{p({\vec{\mathbf{f}}},{\vec{\mathbf{x}}},{\vec{\mathbf{y}}})}{q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})}\right]}_{\triangleq\mathcal{L}}=\underbrace{\operatorname{KL}\!\!\left[q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})\|p({\vec{\mathbf{x}}},{\vec{\mathbf{f}}}|{\vec{\mathbf{y}}})\right]}_{\geq 0}$ (7)
Detailed derivations can be found in Supplement D-A. Maximizing $\mathcal{L}$ w.r.t. $\bm{\theta}$ fine-tunes the model parameters to fit the observed data in the model learning process; while maximizing $\mathcal{L}$ w.r.t. the variational distribution $q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})$ is equivalent to minimizing the KL divergence, $\operatorname{KL}[q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})\|p({\vec{\mathbf{x}}},{\vec{\mathbf{f}}}|{\vec{\mathbf{y}}})]$. That is, it enhances the quality of the variational distribution approximation, bringing it closer to the underlying posterior distribution, $p({\vec{\mathbf{x}}},{\vec{\mathbf{f}}}|{\vec{\mathbf{y}}})$, which corresponds to the model inference process [37]. The capacity of $q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})$ to approximate $p({\vec{\mathbf{x}}},{\vec{\mathbf{f}}}|{\vec{\mathbf{y}}})$ is thus of crucial importance [38].
Based on the model defined in Eq. (6), a generic factorization of $q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})$ is as follows:
$q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})=q({\mathbf{x}}_{0})q({\vec{\mathbf{f}}})\prod_{t=1}^{T}q({\mathbf{x}}_{t}|{\mathbf{f}}_{t}),$ (8)
where $q({\vec{\mathbf{f}}})$ represents the variational distribution of the GP and $q({\mathbf{x}}_{0})\prod_{t=1}^{T}q({\mathbf{x}}_{t}|{\mathbf{f}}_{t})$ corresponds to the variational distribution of the latent states [39, 35]. It is noteworthy that the factorization of the variational distribution presented in Eq. (8) is recognized as an NMF approximation within the GPSSM literature (see formal definition in Appendix C) [28], because it explicitly establishes the dependencies between the latent states and the GP transition function values, as manifested by the terms $\prod_{t=1}^{T}q({\mathbf{x}}_{t}|{\mathbf{f}}_{t})$.
The works most closely related to this paper are probably [27] and [28]. In [27], variational distribution $q({\mathbf{x}}_{t}|{\mathbf{f}}_{t})$ is subjectively set equal to the prior distribution $p({\mathbf{x}}_{t}|{\mathbf{f}}_{t})$, i.e., $q({\mathbf{x}}_{t}|{\mathbf{f}}_{t})=p({\mathbf{x}}_{t}|{\mathbf{f}}_{t})$. But using the prior distribution as an approximation results in no filtering or smoothing effect on latent states, although there is a dependence between the latent states ${\vec{\mathbf{x}}}$ and the transition function values ${\vec{\mathbf{f}}}$. Instead, the work in [28] assumes a parametric Markov-structured Gaussian variational distribution over the temporal states, i.e.,
$q({\mathbf{x}}_{t}|{\mathbf{f}}_{t})={\cal N}({\mathbf{x}}_{t}|\bm{A}_{t}{\mathbf{f}}_{t}+\bm{b}_{t},\bm{S}_{t}),$ (9)
where $\{\bm{A}_{t},\bm{b}_{t},\bm{S}_{t}\}_{t=1}^{T}$ are free variational parameters. This choice, however, introduces a significant drawback: the number of variational parameters grows linearly with the length of the time series. Although this issue can be partially addressed by incorporating an inference network, such as a bidirectional recurrent neural network [26, 39], to learn the variational parameters and make the variational distribution learning with a constant model complexity, fine-tuning these variational parameters of the inference network still requires substantial effort. Consequently, despite the flexibility of the variational distribution described in Eq. (9) allowing for approximation of the true posterior distribution, its empirical performance often falls short of its theoretical expressive capacity [30]. Moreover, it is noteworthy that the inference network commonly takes the input of the observations, ${\vec{\mathbf{y}}}$, rendering the existing inference network-based methods predominantly trained in an offline manner, thus posing challenges when attempting to adapt them for online applications. To address these limitations, we step away from the heavy parameterization (e.g., the black-box inference networks) and propose a novel interpretable EnKF-aided algorithm for variational inference in GPSSMs. Our algorithm can exploit the dependencies between ${\vec{\mathbf{x}}}$ and ${\vec{\mathbf{f}}}$ and alleviate the model and computational complexities while being easily extended to an online learning algorithm.
IV EnVI: EnKF-Aided Variational Inference
This section presents our novel variational inference algorithm for GPSSMs. We begin by introducing the model-based EnKF in Section IV-A, which serves as the foundation for our algorithm. Following that, Section IV-B details the proposed EnKF-aided variational inference algorithm. Lastly, extensive discussions about the properties of our proposed algorithm are given in Section IV-C.
IV-A Ensemble Kalman Filter (EnKF)
The EnKF is a Monte Carlo-based method that excels in handling nonlinear systems compared to the KF and EKF [40]. Given a SSM, see Eq. (II-B), the EnKF sequentially approximates the filtering distributions using $N\!\in\!\mathbb{N}$ equally weighted particles [40]. Specifically, at the prediction step, EnKF first samples particles, ${\mathbf{x}}_{t-1}^{1:N}\triangleq\{{\mathbf{x}}_{t-1}^{n}\}_{n=1}^{N}$, from the filtering distribution, $p({\mathbf{x}}_{t-1}|{\mathbf{y}}_{1:t-1})$, at time $t-1$. Then, the sampled particles are propagated using the state transition function $f(\cdot)$, i.e.,
$\bar{{\mathbf{x}}}_{t}^{n}=f({\mathbf{x}}_{t-1}^{n})+\mathbf{v}_{t}^{n},\ \mathbf{v}_{t}^{n}\sim{\cal N}(0,\mathbf{Q}),\ \forall n.$ (10)
The prediction distribution is then approximated by a Gaussian distribution, i.e. $p({\mathbf{x}}_{t}|{\mathbf{y}}_{1:t-1})\approx{\cal N}({\mathbf{x}}_{t}|\bar{\mathbf{m}}_{t},\bar{\mathbf{P}}_{t})$, where
$\bar{\mathbf{m}}_{t}$ $=\frac{1}{N}\sum_{n=1}^{N}\bar{{\mathbf{x}}}_{t}^{n},$ (11a)
$\bar{\mathbf{P}}_{t}$ $=\frac{1}{N\!-\!1}\sum_{n=1}^{N}(\bar{{\mathbf{x}}}_{t}^{n}-\bar{\mathbf{m}}_{t})(\bar{{\mathbf{x}}}_{t}^{n}-\bar{\mathbf{m}}_{t})^{\top}.$ (11b)
With the linear Gaussian emission in Eq. (II-B), the joint distribution of ${\mathbf{x}}_{t}$ and ${\mathbf{y}}_{t}$ can be readily obtained as follows:
$\begin{bmatrix}{\mathbf{x}}_{t}|{\mathbf{y}}_{1:t-1}\\ {\mathbf{y}}_{t}|{\mathbf{y}}_{1:t-1}\end{bmatrix}\sim{\cal N}\left(\begin{bmatrix}\bar{\mathbf{m}}_{t}\\ \bm{C}\bar{\mathbf{m}}_{t}\end{bmatrix},\begin{bmatrix}\bar{\mathbf{P}}_{t},&\bar{\mathbf{P}}_{t}\bm{C}^{\top}\\ \bm{C}\bar{\mathbf{P}}_{t},&\bm{C}\bar{\mathbf{P}}_{t}\bm{C}^{\top}+\mathbf{R}\end{bmatrix}\right).$ (12)
Thus, the filtering distribution, $p({\mathbf{x}}_{t}|{\mathbf{y}}_{1:t})\!=\!{\cal N}({\mathbf{x}}_{t}|\mathbf{m}_{t},\mathbf{P}_{t}),$ can be obtained using the conditional Gaussian identity at the update step, where
$\mathbf{m}_{t}$ $=\bar{\mathbf{m}}_{t}+\bar{\mathbf{G}}_{t}({\mathbf{y}}_{t}-\bm{C}\bar{\mathbf{m}}_{t}),$ (13a)
$\mathbf{P}_{t}$ $=\bar{\mathbf{P}}_{t}-\bar{\mathbf{P}}_{t}\bm{C}^{\top}\bar{\mathbf{G}}_{t}^{\top},$ (13b)
and $\bar{\mathbf{G}}_{t}=\bar{\mathbf{P}}_{t}\bm{C}^{\top}(\bm{C}\bar{\mathbf{P}}_{t}\bm{C}^{\top}+\mathbf{R})^{-1}$ is the Kalman gain [40]. Each filtered particle ${{\mathbf{x}}}_{t}^{n}$ is then obtained from a Kalman-type update, i.e.,
${{\mathbf{x}}}_{t}^{n}=\bar{{\mathbf{x}}}_{t}^{n}+\bar{\mathbf{G}}_{t}({\mathbf{y}}_{t}+\mathbf{e}_{t}^{n}-\bm{C}\bar{{\mathbf{x}}}_{t}^{n}),\ \mathbf{e}_{t}^{n}\sim{\cal N}(0,\mathbf{R}),\ \forall n.$ (14)
Remark 1.
Compared to another Monte Carlo-based method, PF, EnKF demonstrates computational efficiency, particularly in exceedingly high-dimensional spaces [41], though the PF may offer more flexibility in addressing highly non-Gaussian and nonlinear aspects. More importantly, the inherent differentiable nature of the EnKF (see Eqs. (10)–(14)) enables seamless integration with off-the-shelf automatic differentiation tools (e.g., PyTorch [42]) for model parameter optimization (see more discussions in Section IV-C). This stands in contrast to PF, where the discrete distribution resampling significantly poses challenges for the utilization of the reparameterization trick [43], resulting in significantly higher computational complexity for the parameters gradient computations [44, 45, 46, 47].
In the following subsections, we will describe our novel EnKF-aided NMF variational inference method, which significantly reduces the number of variational parameters and ultimately enhances the learning and inference performance of GPSSMs.
IV-B EnKF-Aided Variational Inference (EnVI)
IV-B1 Sparse GPSSMs
Before introducing the utilization of EnKF in the variational inference framework, let us first introduce the sparse GP [32], a widely utilized technique in various GP variational approximations. This method serves as a scalable approach to model the corresponding GP component within the approximate posterior distribution (see Eq. (8)), thereby guaranteeing inherent scalability in GPSSM. The main idea of the sparse GP is to introduce a small set of inducing points $\vec{{\mathbf{z}}}\triangleq\{{\mathbf{z}}_{i}\}_{i=1}^{M}$ and $\vec{{\mathbf{u}}}\triangleq\{{\mathbf{u}}_{i}\}_{i=1}^{M}$, $M\!\ll\!T$, to serve as the surrogate of the associated GP, where the inducing inputs, ${\mathbf{z}}_{i}\!\in\!\mathbb{R}^{d_{x}},\forall i$, are placed in the same space as the latent states ${\mathbf{x}}_{t}$, while the corresponding inducing outputs ${\mathbf{u}}_{i}\!=\!f({\mathbf{z}}_{i})$ follow the same GP prior as ${\vec{\mathbf{f}}}$. With the augmentation of the inducing points, the joint distribution of the GPSSM becomes
$p({\vec{\mathbf{f}}},{\vec{\mathbf{u}}},{\vec{\mathbf{x}}},{\vec{\mathbf{y}}})=p({\mathbf{x}}_{0})p({\vec{\mathbf{f}}},{\vec{\mathbf{u}}})\prod_{t=1}^{T}p({\mathbf{y}}_{t}|{\mathbf{x}}_{t})p({\mathbf{x}}_{t}|{\mathbf{f}}_{t}),$ (15)
where $p({\vec{\mathbf{f}}},{\vec{\mathbf{u}}})=p({\vec{\mathbf{u}}})p({\vec{\mathbf{f}}}|{\vec{\mathbf{u}}})=p({\vec{\mathbf{u}}})\prod_{t=1}^{T}p({\mathbf{f}}_{t}|{\mathbf{x}}_{t-1},{\vec{\mathbf{u}}})$ is the augmented GP prior and $p({\mathbf{f}}_{t}|{\mathbf{x}}_{t-1},{\vec{\mathbf{u}}})$ is the noiseless GP prediction whose mean and covariance can be computed similarly to Eq. (II-A). The introduced inducing inputs ${\vec{\mathbf{z}}}$ will be treated as variational parameters and jointly optimized with model parameters [48]. We will further describe this later.
Suppose that the inducing outputs ${\vec{\mathbf{u}}}$ serve as sufficient statistics for the GP function values ${\vec{\mathbf{f}}}$, such that given ${\vec{\mathbf{u}}}$, the GP function values ${\vec{\mathbf{f}}}$ and any novel ${\mathbf{f}}_{*}$ are independent [32], i.e., $p({\mathbf{f}}_{*}|{\vec{\mathbf{f}}},{\vec{\mathbf{u}}})=p({\mathbf{f}}_{*}|{\vec{\mathbf{u}}})$ for any ${\mathbf{f}}_{*}$. We can integrate out ${\vec{\mathbf{f}}}$ in Eq. (15), and the transition function is fully characterized using only the inducing points. Consequently, we have:
$p({\vec{\mathbf{u}}},{\vec{\mathbf{x}}},{\vec{\mathbf{y}}})=p({\mathbf{x}}_{0})p({\vec{\mathbf{u}}})\prod_{t=1}^{T}p({\mathbf{y}}_{t}|{\mathbf{x}}_{t})p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1}),$ (16)
where
$p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})$ $=\int p({\mathbf{x}}_{t}|{\mathbf{f}}_{t})p({\mathbf{f}}_{t}|{\mathbf{x}}_{t-1},{\vec{\mathbf{u}}},{\vec{\mathbf{z}}})\mathrm{d}{\mathbf{f}}_{t}$ (17)
$={\cal N}({\mathbf{x}}_{t}~{}|~{}\bm{\xi}_{t},\ \bm{\Xi}_{t}),$
and with a bit abuse of notation,
$\bm{\xi}_{t}=\bm{K}_{{\mathbf{x}}_{t-1},{\vec{\mathbf{z}}}}(\bm{K}_{{\vec{\mathbf{z}}},{\vec{\mathbf{z}}}}+\mathbf{Q})^{\!-\!1}{\vec{\mathbf{u}}},$ (18a)
$\bm{\Xi}_{t}=\bm{K}_{{\mathbf{x}}_{t\!-\!1},{\mathbf{x}}_{t\!-\!1}}\!+\!\mathbf{Q}\!-\!\bm{K}_{{\mathbf{x}}_{t\!-\!1},{\vec{\mathbf{z}}}}(\bm{K}_{{\vec{\mathbf{z}}},{\vec{\mathbf{z}}}}+\mathbf{Q})^{\!-\!1}\bm{K}_{{\mathbf{x}}_{t\!-\!1},{\vec{\mathbf{z}}}}^{\top}.$ (18b)
That is to say, with the aid of sparse GPs, the computational complexity of GPSSMs can be reduced to $\mathcal{O}(d_{x}TM^{2})$, comparing to the original $\mathcal{O}(d_{x}T^{3})$ [35].
IV-B2 ELBO for sparse GPSSM
In the context of the GPSSM described in Eq. (16), we first assume a generic variational distribution for the latent variables, $\{{\vec{\mathbf{u}}},{\vec{\mathbf{x}}}\}$, factorized as follows:
$q({\vec{\mathbf{u}}},{\vec{\mathbf{x}}})$ $=q({\vec{\mathbf{u}}})q({\mathbf{x}}_{0})\prod_{t=1}^{T}\int q({\mathbf{x}}_{t}|{\mathbf{f}}_{t})p({\mathbf{f}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})\mathrm{d}{\mathbf{f}}_{t}$ (19)
$=q({\vec{\mathbf{u}}})q({\mathbf{x}}_{0})\prod_{t=1}^{T}q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1}).$
Here, the variational distribution over the inducing outputs, $q({\vec{\mathbf{u}}})$, is explicitly assumed to be a free-form Gaussian, i.e.,
$q({\vec{\mathbf{u}}})=\prod_{d=1}^{d_{x}}{\cal N}(\{{\mathbf{u}}_{i,d}\}_{i=1}^{M}|{\mathbf{m}}_{d},\mathbf{L}_{d}\mathbf{L}_{d}^{\top})={\cal N}({\vec{\mathbf{u}}}~{}|~{}\mathbf{m},\mathbf{S}),$ (20)
where $\mathbf{m}\!\!=\!\![{\mathbf{m}}_{1}^{\top},\ldots,{\mathbf{m}}_{d_{x}}^{\top}]^{\top}\in\mathbb{R}^{Md_{x}}$ and $\mathbf{S}\!\!=\!\!\operatorname{diag}(\mathbf{L}_{1}\mathbf{L}_{1}^{\top},\ldots,\mathbf{L}_{d_{x}}\mathbf{L}_{d_{x}}^{\top})\in\mathbb{R}^{Md_{x}\times Md_{x}}$ are free variational parameters. This explicit representation of the variational distribution enables scalability through the utilization of stochastic gradient-based optimization [49], as it allows for the independence of individual GP predictions given the explicit inducing points [48]. Similarly, we assume that the variational distribution over the initial state is $q({\mathbf{x}}_{0})=\mathcal{N}({\mathbf{x}}_{0}|{\mathbf{m}}_{0},\mathbf{L}_{0}\mathbf{L}_{0}^{\top})$, where ${{\mathbf{m}}}_{0}\in\mathbb{R}^{d_{x}}$ and lower-triangular matrix $\mathbf{L}_{0}\in\mathbb{R}^{d_{x}\times d_{x}}$ are free variational parameters of $q({\mathbf{x}}_{0})$. The variational distribution of the latent states, $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})$, obtained by integrating out ${\mathbf{f}}_{t}$ as shown in Eq. (19), will be implicitly modeled by resorting to the EnKF technique. This modeling approach will help eliminate the need to parameterize $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})$, a common requirement in the existing works [26, 27, 28, 29], thus overcoming the challenges associated with optimizing a large number of variational parameters, as discussed in Section III.
Before elucidating the methodology of employing EnKF to eliminate the parameterization for the variational distribution $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})$, we first undertake the derivation of the ELBO, which is succinctly summarized as follows.
Theorem 1.
Upon the augmentation of induced points into the GPSSM (see Eq. (16)) and under the NMF assumption of the variational distribution (see Eq. (19)), the model evidence lower bound for joint learning and inference is:
$\!\!\mathcal{L}$ $=\underbrace{\mathbb{E}_{q({\vec{\mathbf{u}}},{\vec{\mathbf{x}}})}\left[\sum_{t=1}^{T}\log p({\mathbf{y}}_{t}|{\mathbf{x}}_{t})\right]}_{\mathrm{term~{}1}}\ -\ \underbrace{\operatorname{KL}[q({\mathbf{x}}_{0})\|p({\mathbf{x}}_{0})]}_{\mathrm{term~{}2}}$ (21)
$\quad\!-\!\underbrace{\operatorname{KL}[q({\vec{\mathbf{u}}})\|p({\vec{\mathbf{u}}})]}_{\mathrm{term~{}3}}-\underbrace{\mathbb{E}_{q({\vec{\mathbf{u}}},{\vec{\mathbf{x}}})}\left[\sum_{t=1}^{T}\log\frac{q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})}{p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})}\right]}_{\mathrm{term~{}4}}.$
Proof.
The proof can be found in Appendix A
An important observation lies in the interpretability of each component within the ELBO in Eq. (21). Specifically, when maximizing the ELBO for learning and inference:
  • Term 1 corresponds to the data reconstruction error, which encourages any state trajectory ${\vec{\mathbf{x}}}$ sampled from the variational distribution, $q({\vec{\mathbf{u}}},{\vec{\mathbf{x}}})$, to accurately reconstruct the observed data.
  • Terms 2 and 3 serve as regularization terms for $q({\mathbf{x}}_{0})$ and $q({\vec{\mathbf{u}}})$, respectively. They discourage significant deviations of the variational distributions from the corresponding prior distributions.
  • Term 4 represents a regularization for $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})$, which discourages significant deviations of $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})$ from the prior $p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})$.
IV-B3 EnVI algorithm
The NMF assumption applied to the variational distribution (Eq. (19)) typically results in an intractable evaluation of the first and fourth terms within the ELBO [28]. We next proceed to outline the approximate evaluation of the ELBO using EnKF. The main idea revolves around utilizing the filtering distribution acquired from EnKF as a proxy for the corresponding variational distribution, acting as an approximation for the underlying posterior distribution, $p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\vec{\mathbf{y}}})$.
Building upon the EnKF outlined in Section IV-A and assuming that we have acquired the variational distribution, $q({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}})\triangleq p({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})\!=\!{\cal N}({\mathbf{x}}_{t-1}|\mathbf{m}_{t-1},\mathbf{P}_{t-1})$, at time step $t\!-\!1$, we can employ the GP transition, presented in Eq. (17), to perform the prediction step and generate $N$ predicted samples $\bar{{\mathbf{x}}}_{t}^{1:N}$, i.e.,
$\bar{{\mathbf{x}}}_{t}^{n}\sim p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},\mathbf{x}_{t-1}^{n}),\ \ n=1,2,\ldots,N,$ (22)
where $\{{\mathbf{x}}_{t-1}^{n}\}_{n=1}^{N}$ are $N$ particles generated from $q({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}})$ with equal weights. With the predicted samples $\bar{{\mathbf{x}}}_{t}^{1:N}$, we thus can approximate the prediction distribution as a Gaussian:
$p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})$ $=\int p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})p({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})\mathrm{d}{\mathbf{x}}_{t-1}$
$=\int p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})q({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}})\mathrm{d}{\mathbf{x}}_{t-1}$
$\approx{\cal N}({\mathbf{x}}_{t}|\bar{\mathbf{m}}_{t},\bar{\mathbf{P}}_{t})$ (23)
where $\bar{\mathbf{m}}_{t}$ and $\bar{\mathbf{P}}_{t}$ can be computed using Eq. (IV-A). Similarly, during the update step, utilizing Eqs. (IV-A) and (14), we can derive the variational distribution (approximate posterior distribution) at time step $t$,
$q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}})\triangleq p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t})={\cal N}({\mathbf{x}}_{t}|\mathbf{m}_{t},\mathbf{P}_{t})$ (24)
and obtain the set of $N$ updated samples ${\mathbf{x}}_{t}^{1:N}$. Note that in the context of GPSSMs, both the prediction and update steps are conditioned on the inducing points, which act as a surrogate for the transition function.
So far, we have introduced the utilization of EnKF to obtain the approximate posterior distribution, $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}})$. However, in order to evaluate the ELBO and compute term 1 and term 4 in Eq. (21), we still need to determine the conditional distribution, $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})$. Despite having that
$q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}})=\int q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})q({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}})\mathrm{d}{\mathbf{x}}_{t-1},$ (25)
and both approximations $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}})$ and $q({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}})$ being Gaussian, determining the functional form of a common posterior transition $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})=p({\mathbf{x}}_{t}|{\mathbf{x}}_{t-1},{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t}),\forall t$, remains a challenging task. Therefore, alternative approximation methods are required to efficiently evaluate the ELBO. One alternative manner is to simply let $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})=p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})$, we then end up with the same ELBO as in [27], where no filtering or smoothing effect is feasible. To circumvent this limitation, we examine the difference between term 1 and term 4 in Eq. (21) and propose our EnKF-aided variational lower bound approximation. The main result is summarized in the following proposition.
Proposition 1.
Under the approximations of EnKF filtering (Eq. (24)) and $p({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t})\approx p({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})$, the ELBO presented in Theorem 1 can be reformulated as a summation over several simple terms:
$\!\!\mathcal{L}\!\approx\!\mathbb{E}_{q({\vec{\mathbf{u}}})}\!\left[\sum_{t=1}^{T}\log p({\mathbf{y}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})\right]$ $\!-\operatorname{KL}[q({\mathbf{x}}_{0})\|p({\mathbf{x}}_{0})]$ (26)
$\!-\operatorname{KL}[q({\vec{\mathbf{u}}})\|p({\vec{\mathbf{u}}})],$
where the log-likelihood, $\log p({\mathbf{y}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})$ in the first term can be analytically evaluated, i.e.,
$\log p({\mathbf{y}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})$ $\!=\!\log\int p({\mathbf{y}}_{t}|{\mathbf{x}}_{t})p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})\mathrm{d}{\mathbf{x}}_{t}$ (27)
$\!=\!\log{\cal N}({\mathbf{y}}_{t}|\bm{C}\bar{\mathbf{m}}_{t},\ \bm{C}\bar{\mathbf{P}}_{t}\bm{C}^{\top}),$
and the two KL divergence terms can also be analytically computed in closed form, due to the Gaussian nature of the prior and variational distributions [10].
Proof.
The proof can be found in Appendix A
Based on the result in Proposition 1, we first utilize the reparametrization trick [43] to sample ${\vec{\mathbf{u}}}\!\sim\!q({\vec{\mathbf{u}}})$ and numerically get an unbiased evaluation of the expected log-likelihood, $\mathbb{E}_{q({\vec{\mathbf{u}}})}\!\left[\sum_{t=1}^{T}\log p({\mathbf{y}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})\right]$; then we apply gradient-based methods [50] to maximize the $\mathcal{L}$ in Eq. (26) w.r.t. the model parameters ${{\bm{\theta}}}=[\bm{\theta}_{gp},\mathbf{Q},\mathbf{R}]$ and the variational parameters $\bm{\zeta}=[\mathbf{m}_{0},\mathbf{L}_{0},\mathbf{m},\mathbf{S},{\vec{\mathbf{z}}}]$. Detailed routine for implementing the EnKF-aided variational learning and inference algorithm, termed as EnVI, is summarized in Algorithm 1.
Algorithm 1 EnKF-aided variational learning and inference
1:Input: ${{\bm{\theta}}}=[{{\bm{\theta}}}_{gp},\mathbf{Q},\mathbf{R}],~{}{\bm{\zeta}},~{}{\mathbf{y}}_{1:T}$, ${\mathbf{x}}_{0}^{1:N}\sim q(x_{0})$
2:while iterations not terminated do
3:     ${\vec{\mathbf{u}}}\sim q({\vec{\mathbf{u}}})$, $L_{\ell}=0$
4:     for $t=1,2,\ldots,T$ do
5:         Get prediction samples using Eq. (22)
6:         Get empirical moments $\bar{\mathbf{m}}_{t},\bar{\mathbf{P}}_{t}$ using Eq. (23)
7:         Get Kalman gain: $\bar{\mathbf{G}}_{t}\!=\!\bar{\mathbf{P}}_{t}\bm{C}^{\top}(\bm{C}\bar{\mathbf{P}}_{t}\bm{C}^{\top}\!\!+\!\mathbf{R})^{-1}$
8:         Get updated samples using Eq. (14)
9:         Evaluate the log-likelihood using Eq. (27), and
$L_{\ell}=L_{\ell}+\log p({\mathbf{y}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})$
10:     end for
11:     $\mathcal{L}=L_{\ell}-\operatorname{KL}(q({\mathbf{x}}_{0})\|p({\mathbf{x}}_{0}))-\operatorname{KL}(q({\vec{\mathbf{u}}})\|p({\vec{\mathbf{u}}}))$
12:     Maximize $\mathcal{L}$ and update ${{\bm{\theta}}}$, ${\bm{\zeta}}$ using Adam [50]
13:end while
14:Output: EnKF particles ${\mathbf{x}}_{0:T}^{1:N}$, model parameters ${{\bm{\theta}}}$, and variational parameters ${\bm{\zeta}}$.
It is noteworthy that the newly derived ELBO, $\mathcal{L}$ circumvents the explicit evaluation of the first and fourth terms in Eq. (21), and sidesteps the computational challenges posed by the heavy parameterization of $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})$. Consequently, EnVI can substantially improve the efficiency of model learning and inference. In addition, maximizing the ELBO in Eq. (26) can be interpreted as follows: The objective is to optimize the model parameters and variational parameters such that the GPSSM can fit the observed data well at each step (indicated by the first term); simultaneously, the KL regularization terms impose constraints to prevent model overfitting (indicated by the second and third terms).
Remark 2.
The computational complexity of the EnVI algorithm predominantly lies in the evaluation of $N$ independent particles on the GP transition during the prediction step (see Eq. (22)). Recall that the number of inducing points is significantly smaller than the length of the data sequence, i.e., $M\!\ll\!T$, and assume that $M\!\geq\!d_{x}$. In this context, the computational complexity of Algorithm 1 scales as $\mathcal{O}(NTd_{x}M^{2})$. In practice, $N$ is often a small number, and the computation of the $N$ particles on the GP transition can be run in parallel [42], resulting in the computational complexity in real-world deployments scaling as $\mathcal{O}(Td_{x}M^{2})$. This cost matches that of existing works. Yet, the streamlined model complexity in EnVI enhances its computational robustness and accelerates convergence compared to the existing works.
IV-C More Discussions and Insights
This subsection presents more detailed insights into the proposed EnVI and discusses its connections to existing works.
First of all, the importance of the differentiable nature within EnKF, as mentioned in Remark 1, to the EnVI becomes more apparent. The inherent differentiability in EnVI, spanning from parameters ($\{{{\bm{\theta}}},{\bm{\zeta}}\}$) to latent states and extending to the objective function (the ELBO), enables principled joint learning and inference using modern off-the-shelf automatic differentiation tools (e.g., PyTorch [42]). This contrasts with most existing EnKF-based dynamical systems learning methods (see, e.g., [51, 52] and the references therein), which employ the expectation-maximization (EM) algorithm [10] to iteratively update the model parameters ${{\bm{\theta}}}$ and latent state trajectory ${\vec{\mathbf{x}}}$. It has been reported that such EM-based methods disregard the gradient information from the parameters ${{\bm{\theta}}}$ to the state trajectory ${\vec{\mathbf{x}}}$, which can potentially degrade the learning performance [47, 53]. In contrast, by regarding the latent state trajectory as a function of both the model parameters ${{\bm{\theta}}}$ and variational parameters ${\bm{\zeta}}$, our method can jointly optimize these parameters in a principled way, resulting in enhanced performance.
Second, it is worth noting that the proposed EnVI algorithm falls under the NMF category (see definition in Appendix C), given the exploitation of the dependencies between latent states and the transition function, as evident in the variational distribution, $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}})$ in Eq. (24). This indicates that EnVI inherits the favorable characteristics of NMF algorithms [27], which have the potential to enhance learning accuracy and address the issue of underestimating state inference uncertainty, commonly encountered in MF algorithms [33].
Last but not least, our method offers an effective means of mitigating the risk of overfitting. This is mainly due to the fact that our method leverages the Bayesian nonparametric model and the variational inference framework to derive the ELBO, as presented in Eq. (26), from which the additional regularization terms for the initial state and state transition function can effectively mitigate the overfitting issue. As a comparison, in the state-of-the-art EnKF-based dynamical system learning method, autodifferentiable EnKF (AD-EnKF) [47, 51], the transition function in the SSM is modeled using a deterministic parametric model, specifically a neural network, and the optimization objective function is the logarithm of marginal likelihood of the model, i.e.
$\mathcal{L}_{\text{AD-EnKF}}=\log p({\mathbf{y}}_{1:T})=\sum_{t=1}^{T}\log p({\mathbf{y}}_{t}|{\mathbf{y}}_{1:t-1}).$ (28)
Maximizing this objective function solely w.r.t. the model parameters can easily lead to overfitting. We defer further discussions of this issue to Section VI.
V OEnVI: Online Implementation of EnVI
In this section, we further explore online setting where data is processed sequentially, one sample at a time. It is within this context that the simultaneous inference of states and nonlinear dynamics in GPSSMs presents significant challenges [54, 55, 56]. The good news is that our EnVI approach readily lends itself to online learning scenarios. At each time step $t$, we can naturally maximize the corresponding objective function, denoted as $L_{\ell_{t}}$, given by
$L_{\ell_{t}}=\mathbb{E}_{q({\vec{\mathbf{u}}})}\left[\log p({\mathbf{y}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})\right]-\operatorname{KL}(q({\vec{\mathbf{u}}})\|p({\vec{\mathbf{u}}})),$ (29)
in terms of both the model parameters and variational parameters. Detailed steps for implementing the online EnVI, termed OEnVI, is summarized in Algorithm 2.
Algorithm 2 OEnKF: Online EnVI (Step $t$)
1:Input: ${{\bm{\theta}}}=[{{\bm{\theta}}}_{gp},\mathbf{Q},\mathbf{R}],~{}{\bm{\zeta}},~{}{\mathbf{y}}_{t}$, ${\mathbf{x}}_{t-1}^{1:N}$
2:while iterations not terminated do
3:     ${\vec{\mathbf{u}}}\sim q({\vec{\mathbf{u}}})$, $L_{\ell_{t}}=0$
4:     Get prediction samples using Eq. (22)
5:     Get empirical moments $\bar{\mathbf{m}}_{t},\bar{\mathbf{P}}_{t}$ using Eq. (23)
6:     Get Kalman gain: $\bar{\mathbf{G}}_{t}\!=\!\bar{\mathbf{P}}_{t}\bm{C}^{\top}(\bm{C}\bar{\mathbf{P}}_{t}\bm{C}^{\top}\!\!+\!\mathbf{R})^{-1}$
7:     Get updated samples using Eq. (14)
8:     Evaluate the objective function
$L_{\ell_{t}}=\log p({\mathbf{y}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})-\operatorname{KL}(q({\vec{\mathbf{u}}})\|p({\vec{\mathbf{u}}}))$
9:     Maximize $L_{\ell_{t}}$ and update ${{\bm{\theta}}}$, ${\bm{\zeta}}$ using Adam [50]
10:end while
11:Output: EnKF particles ${\mathbf{x}}_{t}^{1:N}$, model parameters ${{\bm{\theta}}}$, and variational parameters ${\bm{\zeta}}$.
Remark 3.
Analogous to EnVI, the computational complexity of OEnVI scales as $\mathcal{O}(d_{x}M^{2})$ under practical parallel computing environments.
An interesting insight for maximizing the objective function of OEnVI in Eq. (29) (as well as the EnVI’s objective function in Eq. (26)) is that it essentially encourages successful data reconstruction while simultaneously ensuring that the filtering distribution $p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t})$ and the prediction distribution $p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})$ do not deviate too far from each other, apart from the regularization of $q({\vec{\mathbf{u}}})$. This result is supported by the following proposition.
Proposition 2.
The log-likelihood, $\log p({\mathbf{y}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})$, essentially is the difference between the data reconstruction error, represented by $\mathbb{E}_{q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}})}\left[\log p({\mathbf{y}}_{t}|{\mathbf{x}}_{t})\right]$, and the KL divergence between the variational distribution $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t})$ and the prediction distribution $p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})$. That is,
$\log p({\mathbf{y}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})$ (30)
$\!=\!\mathbb{E}_{q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}})}[\log p({\mathbf{y}}_{t}|{\mathbf{x}}_{t})]\!-\!\operatorname{KL}\left[q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}})\|p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})\right].$
Thus, an alternative objective function for OEnVI can be expressed as
$L_{\ell_{t}}$ $=\mathbb{E}_{q({\vec{\mathbf{u}}})}\left[\mathbb{E}_{q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}})}[\log p({\mathbf{y}}_{t}|{\mathbf{x}}_{t})]\right]$ (31)
$\quad-\mathbb{E}_{q({\vec{\mathbf{u}}})}\left[\operatorname{KL}\left[q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}})\|p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})\right]\right]$
$\quad-\operatorname{KL}\left[q({\vec{\mathbf{u}}})\|p({\vec{\mathbf{u}}})\right]$
Proof.
The proof can be found in Appendix B
This insight sheds light on the interplay between data reconstruction and the alignment of filtering and prediction distributions in the EnVI and OEnVI algorithms.
Up to this point, it is worth noting that in contrast to the existing inference network-based variational algorithms [26, 27, 28, 29, 35, 39, 56, 55], OEnVI is a simple and straightforward extension of EnVI, which benefits from eliminating the dependence on the additional parametric variational distributions. Previous works have typically employed inference networks that take the entire sequence of observations ${\mathbf{y}}_{1:T}$ as input, necessitating a significant amount of data for offline training. While it is conceivable to limit the input length of the inference network to a shorter sequence $l$, such as ${\mathbf{y}}_{t-l:t}$, this approach still leads to prolonged training times and higher computational requirements for optimizing the inference network parameters [56]. Moreover, storing historical inputs ${\mathbf{y}}_{t-l:t}$ adds to the storage overhead, making it problematic in situations where historical data duplication and storage are not permissible. In sharp contrast, OEnVI successfully overcomes the aforementioned challenges related to the optimization of inference networks. As a result, it facilitates more efficient learning and inference processes, contributing to potential improved overall performance. Furthermore, OEnVI offers a more principled objective, see Eq. (31), by simultaneously minimizing the KL divergence, accounting for data reconstruction error balance, and applying regularization to the transition function to mitigate model overfitting.
Figure 2: EnVI (top) & OEnVI (bottom) on state inference in linear Gaussian SSM. The RMSE of the latent state estimates for KF, EnVI, and OEnVI are 0.5252, 0.6841, and 0.7784, respectively; the RMSE between the observations and the latent states is 0.9872.
VI Experiments and Results
This section presents a comprehensive numerical study of the proposed EnVI and OEnVI. Section VI-A showcases the filtering performance. In Section VI-B, we present the system dynamics learning performance. The series forecasting performance of EnVI on various real datasets is illustrated in Section VI-C. Finally, Section VI-D provides a comprehensive demonstration of the performance of the OEnVI online algorithm. More details regarding the experimental setup can be found in supplementary material, and the accompanying source code is publicly available online [^2] https://github.com/zhidilin/gpssmProj.
VI-A Learning and Inference in Linear Gaussian SSMs
We investigate the learning and inference capacity of EnVI and OEnVI, by using a linear Gaussian state-space model (LGSSM) where exact inference of latent state is applicable (i.e. KF). Specifically, we use the following LGSSM, a car tracking example given in the textbook [2], to generate observation data, ${\mathbf{y}}_{1:T}$, for training EnVI and OEnVI,
$\mathbf{x}_{t}$ $=\mathbf{H}\mathbf{x}_{t-1}+\mathbf{v}_{t-1},$ $\mathbf{v}_{t-1}$ $\sim\mathcal{N}(\mathbf{0},\mathbf{Q}),$ (32a)
$\mathbf{y}_{t}$ $=\bm{C}\mathbf{x}_{t}+\mathbf{e}_{t},$ $\mathbf{e}_{t}$ $\sim\mathcal{N}(\mathbf{0},\mathbf{R}),$ (32b)
where the state and the observation are both four dimensional, and $\bm{C}=\bm{I}_{4\times 4},\mathbf{R}=\sigma_{\mathrm{R}}^{2}\bm{I}_{4\times 4}$ with $\sigma_{\mathrm{R}}=0.5$;
$\mathbf{H}=\left(\begin{array}[]{cccc}1&0&\Delta t&0\\ 0&1&0&\Delta t\\ 0&0&1&0\\ 0&0&0&1\end{array}\right)$ (33)
and
$\mathbf{Q}=\left(\begin{array}[]{cccc}\frac{q_{1}^{\mathrm{c}}\Delta t^{3}}{3}&0&\frac{q_{1}^{\mathrm{c}}\Delta t^{2}}{2}&0\\ 0&\frac{q_{2}^{\mathrm{c}}\Delta t^{3}}{3}&0&\frac{q_{2}^{\mathrm{c}}\Delta t^{2}}{2}\\ \frac{q_{1}^{\mathrm{c}}\Delta t^{2}}{2}&0&q_{1}^{\mathrm{c}}\Delta t&0\\ 0&\frac{q_{2}^{\mathrm{c}}\Delta t^{2}}{2}&0&q_{2}^{\mathrm{c}}\Delta t\end{array}\right)$ (34)
with $\Delta t=0.1,q_{1}^{\mathrm{c}}=q_{2}^{\mathrm{c}}=1$.
We begin by generating $T=120$ training observations. For EnVI, we employ $1000$ epochs/iterations for training, but convergence is typically achieved approximately $300$ iterations. In OEnVI, the parameters ${{\bm{\theta}}}$ and ${\bm{\zeta}}$ are updated once per time step $t$. Both EnVI and OEnVI employ $15$ inducing points, a setting that will be used for subsequent experiments unless otherwise specified. We report the state inference results, which are depicted in Fig. 2. It can be observed that the state inference performance of EnVI and OEnVI is comparable to that of the KF in terms of state-fitting root mean square error (RMSE), despite being trained solely on noisy observations without any physical model knowledge.
Another finding is that though OEnVI incurs a lower training cost compared to EnVI, this advantage comes at the expense of inadequate learning of the latent dynamics, leading to a less accurate estimation of the latent states when compared to EnVI. The discrepancy is evident in Fig. 2, where OEnVI exhibits larger estimation RMSE and greater estimation uncertainty for the latent states in comparison to EnVI and KF. Notably, EnVI relies on offline training, resulting in an uncertainty quantification that closely approaches the optimal estimate, the KF estimate. Nevertheless, it is essential to mention that, with continuous online data arrival, OEnVI can eventually achieve a comparable state estimation performance as EnVI. We observed that after observing $360$ data points, OEnVI achieves a latent state RMSE estimation of $0.6512$. Further details on this aspect of the results are provided in the supplementary material (See Supplement E-A).
TABLE I: Comparison of our method with other competitors on the kink function dataset. Shown are mean and standard errors over five repetitions of the fitting MSE (lower is better) and the log-density (higher is better) of the kink function varying the emission noise variance $\sigma_{\mathrm{R}}^{2}$.
Method $\sigma_{\mathrm{R}}^{2}=0.008$ (MSE $\mid$ Log-Likelihood) $\sigma_{\mathrm{R}}^{2}=0.08$ (MSE $\mid$ Log-Likelihood) $\sigma_{\mathrm{R}}^{2}=0.8$ (MSE $\mid$ Log-Likelihood)
vGPSSM [26] $1.0410\!\pm\!0.7426\mid\!-\!27.5981\!\pm\!19.7817$ $1.6390\!\pm\!0.6783\mid\!-\!30.9557\!\pm\!16.9218$ $1.9584\!\pm\!0.9655\mid\!-\!56.5997\!\pm\!37.8221$
VCDT [28] $0.2057\!\pm\!0.2219\mid\!-\!1.058\!\pm\!1.5005$ $0.1934\!\pm\!0.0140\mid\!-\!0.5867\!\pm\!0.2610$ $1.4035\!\pm\!0.6470\mid\!-\!3.8092\!\pm\!0.6588$
AD-EnKF [47] $0.0285\!\pm\!0.0318\mid\!-\!3.6282\!\pm\!6.3514$ $1.5246\!\pm\!0.9734\mid\!-\!242.2795\!\pm\!194.6741$ $1.3489\!\pm\!0.3102\mid\!-\!267.7068\!\pm\!62.0488$
EnVI (ours) $\bm{0.0046}\!\pm\!\bm{0.0025}\mid\bm{1.1060}\!\pm\!\bm{0.0381}$ $\bm{0.0536}\!\pm\!\bm{0.0232}\mid\bm{0.1025}\!\pm\!\bm{0.1075}$ $\bm{0.5315}\!\pm\!\bm{0.1542}\mid\bm{\!-\!1.0439}\!\pm\!\bm{0.1714}$
(a) vGPSSM
(b) VCDT
(c) AD-EnKF
(d) EnVI
Figure 3: Kink transition function learning performance (mean $\pm$ $2\sigma$) using various methods across different levels of emission noise ($\sigma_{\mathrm{R}}^{2}\in\{0.008,0.08,0.8\}$, from left to right).
VI-B System Dynamics Learning
This subsection demonstrates the superior capability of EnVI in learning latent dynamics for GPSSMs. To evaluate its performance, we utilize a 1-D synthetic dataset called the kink function dataset, which is generated from a dynamical system described by Eq. (VI-B), where the nonlinear, smooth and time-invariant function $f({\mathbf{x}}_{t})$ is called the “kink” function,
${\mathbf{x}}_{t+1}=\underbrace{0.8+\left({\mathbf{x}}_{t}+0.2\right)\left[1-\frac{5}{1+\exp(-2{\mathbf{x}}_{t})}\right]}_{\triangleq\text{ ``kink function'' }f({\mathbf{x}}_{t})}+\mathbf{v}_{t},$ (35a)
${\mathbf{y}}_{t}={\mathbf{x}}_{t}+\mathbf{e}_{t},\ \mathbf{v}_{t}\sim{\cal N}(0,\sigma_{\mathrm{Q}}^{2}),\ \mathbf{e}_{t}\sim{\cal N}(0,\sigma_{\mathrm{R}}^{2}).$ (35b)
It is worth mentioning that this specific dynamical system has been extensively employed in GPSSM literature to evaluate the accuracy of the learned GP transition posterior [30].
In showcasing the superior performance of EnVI, we compare it against several prominent competing methods, namely vGPSSM [26], VCDT [28], and AD-EnKF [47]. Our implementation of VCDT incorporates an inference network to address the linear growth in the number of variational parameters. The vGPSSM method adheres to the original paper’s implementation [26], while the AD-EnKF is utilized as per the default software package available online [^3] https://github.com/ymchen0/torchEnKF. The training data sequences are generated by fixing $\sigma_{\mathrm{Q}}^{2}$ at 0.05 and systematically vary $\sigma_{\mathrm{R}}^{2}$ within the range of $\{0.008,0.08,0.8\}$. As a result, we generate three sets of $T=600$ observations each for training. To ensure a fair comparison in the latent space, we adhere to Ialongo et al. [28] and keep the emission model fixed to the true generative ones for all methods, while allowing the transition to be learned. Further details, including the description of the setup for the aforementioned algorithms, are provided in Supplement E-B and the accompanying source code. The result is depicted in Table I and visualized in Fig. 3. We can find that EnVI consistently outperforms existing approaches in terms of the latent dynamic learning performance and robustness. We next conduct two ablation studies.
EnVI vs. Inference Network-Based Methods. Based on the numerical results presented in Table I, we can find that the EnVI exhibits superior dynamic learning performance compared to vGPSSM and VCDT, both of which rely on an inference network. Specifically, as illustrated in Fig. 3, vGPSSM faces more challenges in dynamic learning, while VCDT, categorized under the NMF paradigm, only performs well under more minor noise conditions ($\sigma_{\mathrm{R}}^{2}=0.008$ and $\sigma_{\mathrm{R}}^{2}=0.08$). In contrast, EnVI effectively learns the GP transition well, even in the high noise setting with $\sigma_{\mathrm{R}}^{2}=0.8$.
The primary reason for this discrepancy is the increased model and computational complexity arising from additional inference network parameters, which hinder effective training. Furthermore, the inference network-based methods are prone to convergence into various unfavorable local optima, demonstrating reduced robustness. As a consequence, the learning performance of such methods often fluctuates significantly. In contrast, EnVI inherits the benefits of EnKF and avoids the need to optimize additional variational parameters from the inference network, making it more amenable to optimization and demonstrating enhanced robustness. Our experimental findings consistently demonstrate that EnVI exhibits significantly faster convergence compared to vGPSSM and VCDT. For instance, when considering $\sigma_{\mathrm{R}}^{2}=0.008$, EnVI achieves convergence after 300 iterations, while both vGPSSM and VCDT require much more iterations, as shown in Fig. 4.
Figure 4: Kink function learning performance against the training iterations. EnVI exhibits much faster convergence compared to vGPSSM and VCDT.
EnVI vs. AD-EnKF. The primary difference between EnVI and AD-EnKF lies in their data-driven modules. AD-EnKF employs a parametric model, specifically a neural network (so it is known as a deep SSM (DSSM)), while EnVI utilizes a nonparametric GP. Consequently, the dynamics learned by AD-EnKF exhibit a tendency to be over-confident, as depicted in Fig. 3, as it does not account for uncertainties from the learned transition function. Moreover, the absence of regularization during the training of the neural network in AD-EnKF (as indicated in Eq. (28)) renders the method susceptible to overfitting and being trapped in suboptimal solutions during the training process. As a result, the performance of different repetitions can vary significantly, as shown in Table I.
VI-C Time Series Data Forecasting
This subsection demonstrates the series prediction performance of the proposed EnVI algorithm on five public real-world system identification datasets [^4] https://homes.esat.kuleuven.be/~smc/daisy/daisydata.html, which consist of one-dimensional time series of varying lengths between $296$ to $1024$ data points. In addition to the comparison methods discussed in Section VI-B, EnVI is also compared with several other competitors, including two NMF class methods, PRSSM [27], ODGPSSM [36], and two inference network-based methods, DKF [39] and CO-GPSSM [35], as depicted in Table II. For each method, the first half of the sequence in every dataset is utilized as training data, with the remaining portion designated for testing. Standardization of all datasets is conducted based on the training sequence, and the latent state dimension, $d_{x}$, is consistently set to $4$ for all datasets. Table II reports the series prediction results, wherein the RMSE is averaged over 50-step ahead forecasting.
Table II reveals that EnVI outperforms almost all approaches across five datasets. Specifically, EnVI demonstrates superior performance among the MF and NMF methods. Compared to PRSSM and ODGPSSM [27, 36], which assume equality between variational and prior distributions of latent states, EnVI employs EnKF to filter latent states, leading to an enhanced system dynamics learning performance, and consequently, improving the series predictions. Compared to the inference network-based methods, like VCDT [28] and the MF class methods vGPSSM [26] and CO-GPSSM [35], EnVI eliminates the need to optimize inference network parameters. From an optimization solution perspective, optimizing the inference network leads to a vast solution space for the variational distribution. Consequently, despite their adequate approximation capabilities for the true posterior distribution, these inference network-based methods often fall short of realizing their theoretical potential in empirical performance due to numerous bad local optimums [30]. In contrast, EnVI imposes model-based constraints on the variational distribution by the EnKF, narrowing the solution space and yielding significantly improved and robust empirical performance.
Compared to the DSSM methods, EnVI offers performance advantages due to its non-parametric GP model. In contrast to DKF [39], which utilizes neural networks to model variational distributions and nonlinear SSMs, EnVI employs GPs with much less model parameters, making it particularly suitable for small datasets. While AD-EnKF [47] outperforms DKF, its deterministic neural network modeling approach, and absence of regularization in its objective function cause it to lag behind EnVI in forecasting performance.
TABLE II: Prediction performance (RMSE) of the different models on the system identification datasets. Mean and standard deviation of the prediction results are shown across five seeds. The top three results are highlighted in bold.
Category Method Actuator Ball Beam Drive Dryer Gas Furnace
DSSMs DKF [39] $1.204\pm 0.250$ $0.144\pm 0.005$ $0.735\pm 0.001$ $1.465\pm 0.087$ $5.589\pm 0.066$
AD-EnKF [47] $0.705\pm 0.117$ $0.057\pm 0.006$ $0.756\pm 0.114$ $0.182\pm 0.053$ $1.408\pm 0.090$
MF-based Methods vGPSSM [26] $1.640\pm 0.011$ $0.268\pm 0.414$ $0.740\pm 0.010$ $0.822\pm 0.002$ $3.676\pm 0.145$
CO-GPSSM [35] $0.803\pm 0.011$ $0.079\pm 0.018$ $0.736\pm 0.007$ $0.366\pm 0.146$ $1.898\pm 0.157$
NMF-based Methods PRSSM [27] $0.691\pm 0.148$ $0.074\pm 0.010$ $\bm{0.647}\pm\bm{0.057}$ $0.174\pm 0.013$ $1.503\pm 0.196$
ODGPSSM [36] $0.666\pm 0.074$ $0.068\pm 0.006$ $0.708\pm 0.052$ $0.171\pm 0.011$ $1.704\pm 0.560$
VCDT [28] $0.815\pm 0.012$ $0.065\pm 0.005$ $0.735\pm 0.005$ $0.667\pm 0.266$ $2.052\pm 0.163$
EnVI (ours) $\bm{0.657}\pm\bm{0.095}$ $\bm{0.055}\pm\bm{0.002}$ $0.703\pm 0.050$ $\bm{0.125}\pm\bm{0.017}$ $\bm{1.388}\pm\bm{0.123}$
(a) True and inferred latent trajectory using OEnVI
(b) True and inferred latent trajectory using SVMC
(c) True and inferred latent trajectory using VJF
(d) Filtering and prediction using OEnVI (ours)
(e) Filtering and prediction using SVMC [55]
(f) Filtering and prediction using VJF [56]
Figure 5: Online ${}^{\circledR}$ dynamics learning results of the three online methods [57]
VI-D Online Learning and Inference Using OEnVI
We next evaluate the performance of OEnVI. We compare OEnVI with two very recent competitive online learning algorithms, specifically SVMC [55], which models the variational distribution of latent states using sample particles, and VJF [56], where the variational distribution of latent states is modeled by an inference network. More numerical results of OEnVI about learning kink dynamical function can be found in Supplement E-D.
We evaluate the three online learning methods using ${}^{\circledR}$ data, a dataset previously utilized in [55], showcasing dynamics akin to recurrent switching linear dynamical systems [57]. The simulated trajectory faithfully reproduces the layout of a ${}^{\circledR}$ track, as depicted in Fig. 5a. We train these three methods with $2000$ observations and test them with $500$ observations, both simulated from ${\mathbf{y}}_{t}=\bm{C}{\mathbf{x}}_{t}+\mathbf{e}_{t}$ where $\bm{C}$ is a 10-by-2 matrix generated randomly. The SVMC and VJF algorithms are implemented using the code publicly provided online [^5] https://github.com/catniplab/svmc [^6] https://github.com/catniplab/vjf.
Figs. 5a5c depict the true states (in blue) and the latent states (in red) inferred by the three methods. The results clearly indicate that EnVI and SVMC swiftly captured the real state and maintained accuracy, while VJF faced challenges, primarily due to its difficulty in optimizing the parameters in the inference network. The detailed comparison with SVMC, illustrated in Figs. 5d5f, highlights the superior accuracy of OEnVI in both inference and prediction of latent states. These empirical findings emphasize the efficacy of OEnVI, particularly in its use of EnKF to approximate the variational distribution, demonstrating its advancement over VJF using inference networks, and SVMC using PF methods.
VII Conclusion
In this paper, we have introduced EnVI, a novel NMF algorithm tailored for GPSSMs, which integrates EnKF into a variational inference framework. Additionally, we have presented an extended online version, OEnVI. Both algorithms eliminate the necessity for heavy parameterization like inference networks and shape the variational distribution over latent states through the model-based EnKF. Leveraging the inherent differentiable nature along with the modern automatic differentiation tools, the proposed EnVI and OEnVI can enhance efficiency and algorithmic robustness while improving learning and inference performance compared to existing methods. Detailed analysis and fresh insights for the proposed algorithms are provided to enhance their interpretability. Empirical experiments conducted on diverse real and synthetic datasets, evaluating filtering, prediction, and dynamics learning performance, unequivocally support the effectiveness of the proposed methods.
Appendix A Model Evidence Lower Bound Evaluations
With the model joint distribution, Eq. (16), and the variational distribution, Eq. (19), we can write down the ELBO according to the general definition given in Eq. (7), i.e.,
$\mathcal{L}=\mathbb{E}_{q({\vec{\mathbf{u}}},{\vec{\mathbf{x}}})}\left[\log\frac{p({\vec{\mathbf{u}}},{\vec{\mathbf{x}}},{\vec{\mathbf{y}}})}{q({\vec{\mathbf{u}}},{\vec{\mathbf{x}}})}\right]$ (36)
$=\mathbb{E}_{q({\vec{\mathbf{u}}},{\vec{\mathbf{x}}})}\!\left[\log\frac{p(\mathbf{x}_{0})p({\vec{\mathbf{u}}})\prod_{t=1}^{T}p(\mathbf{x}_{t}|{{\vec{\mathbf{u}}}},{\mathbf{x}}_{t-1})p(\mathbf{y}_{t}|\mathbf{x}_{t})}{q({\vec{\mathbf{u}}})q({\mathbf{x}}_{0})\prod_{t=1}^{T}q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})}\right]$
$=\underbrace{\mathbb{E}_{q({\vec{\mathbf{u}}},{\vec{\mathbf{x}}})}\left[\sum_{t=1}^{T}\log p({\mathbf{y}}_{t}|{\mathbf{x}}_{t})\right]}_{\text{term 1}}\!-\underbrace{\operatorname{KL}[q({\mathbf{x}}_{0})\|p({\mathbf{x}}_{0})]}_{\text{term 2}}$
$\quad\!-\!\underbrace{\operatorname{KL}[q({\vec{\mathbf{u}}})\|p({\vec{\mathbf{u}}})]}_{\text{term 3}}-\underbrace{\mathbb{E}_{q({\vec{\mathbf{u}}},{\vec{\mathbf{x}}})}\left[\sum_{t=1}^{T}\log\frac{q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})}{p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})}\right]}_{\text{term 4}}.$
where the KL divergence terms can be analytically computed in closed form [10] due to the Gaussian nature of the prior and variational distributions. The evaluation of the first and fourth terms is typically intractable. We examine the difference between term 1 and term 4 in Eq. (36) and can have the following lemma.
Lemma 1.
Under the approximations that $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}})=p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t})$, and $p({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t})\approx p({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})$, $\forall t$, computing the difference between term 1 and term 4 in the ELBO (Eq. (21)) yields an approximation to an expected log-likelihood, i.e.,
$\operatorname{term~{}1}-\operatorname{term~{}4}\approx\mathbb{E}_{q({\vec{\mathbf{u}}})}\left[\log p({\mathbf{y}}_{1:T}|{\vec{\mathbf{u}}})\right]$ (37)
Proof.
According to the ELBO given in Eq. (36), we have:
[…]
$\!=\!\mathbb{E}_{q({\vec{\mathbf{u}}},{\vec{\mathbf{x}}})}\!\left[\sum_{t=1}^{T}\log\frac{p({\mathbf{y}}_{t}|{\mathbf{x}}_{t})p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})}{q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})}\right]$ (38a)
$\!=\!\mathbb{E}_{q({\vec{\mathbf{u}}},{\vec{\mathbf{x}}})}\!\left[\sum_{t=1}^{T}\log\frac{p({\mathbf{y}}_{t}|{\mathbf{x}}_{t})p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})q({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}})}{q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})q({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}})}\right]$ (38b)
$\!\approx\!\mathbb{E}_{q({\vec{\mathbf{u}}},{\vec{\mathbf{x}}})}\!\left[\sum_{t=1}^{T}\log\frac{p({\mathbf{y}}_{t}|{\mathbf{x}}_{t})p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})q({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}})}{q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})p({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t})}\right]$ (38c)
$\!=\!\mathbb{E}_{q({\vec{\mathbf{u}}},{\vec{\mathbf{x}}})}\left[\sum_{t=1}^{T}\log\frac{p({\mathbf{y}}_{t},{\mathbf{x}}_{t},{\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})}{p({\mathbf{x}}_{t},{\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1},{\mathbf{y}}_{t})}\right]$ (38d)
$\!=\!\mathbb{E}_{q({\vec{\mathbf{u}}})}\left[\sum_{t=1}^{T}\log p({\mathbf{y}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})\right],$ (38e)
where Eq. (38c) holds because of the approximation of $p({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t})\approx p({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})$ and Eq. (38e) is derived straightforwardly by applying Bayes’ theorem. ∎
According to Lemma 1, and the ELBO given in Eq. (36), we immediately get the following approximated ELBO:
$\!\!\mathcal{L}\!\approx\!\mathbb{E}_{q({\vec{\mathbf{u}}})}\!\left[\sum_{t=1}^{T}\log p({\mathbf{y}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})\right]$ $\!-\operatorname{KL}[q({\mathbf{x}}_{0})\|p({\mathbf{x}}_{0})]$ (39)
$\!-\operatorname{KL}[q({\vec{\mathbf{u}}})\|p({\vec{\mathbf{u}}})],$ (40)
where the log-likelihood, $\log p({\mathbf{y}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})$ in Eq. (39) can be analytically evaluated, see Eq. (27), due to the Gaussian prediction distribution, see Eq. (23), and the linear emission model.
Appendix B Proof of Proposition 2
Proof.
Under the approximation that $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}})\triangleq p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t})$, we have the log-likelihood term
$~{}\log p({\mathbf{y}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})$ (41a)
$=\mathbb{E}_{p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t})}\left[\log p({\mathbf{y}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})\right]$ (41b)
$=\mathbb{E}_{p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t})}\left[\log\frac{p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})p({\mathbf{y}}_{t}|{\mathbf{x}}_{t})}{p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1},{\mathbf{y}}_{t})}\right]$ (41c)
$=-\mathbb{E}_{p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t})}\left[\log\frac{p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t})}{p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})p({\mathbf{y}}_{t}|{\mathbf{x}}_{t})}\right]$ (41d)
$=-\operatorname{KL}\left[p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t})\|p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})\right]$
$\quad~{}+\mathbb{E}_{p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t})}\left[\log p({\mathbf{y}}_{t}|{\mathbf{x}}_{t})\right],$ (41e)
which completes the proof. Here Eq. (41c) is obtained straightforwardly by applying Bayes’ theorem. This result sheds light on the interplay between data reconstruction and the alignment of filtering and prediction distributions in the EnVI and OEnVI algorithms. ∎
Appendix C Mean-Field and Non-Mean-Field Approximations
Definition 1.
If the variational distribution, $q({\vec{\mathbf{f}}},{\vec{\mathbf{u}}},{\vec{\mathbf{x}}})$, is factorized such that the transition function values and the latent states are independent, i.e.,
$q({\vec{\mathbf{f}}},{\vec{\mathbf{u}}})q({\mathbf{x}}_{0})\prod_{t=1}^{T}q({\mathbf{x}}_{t}|{\mathbf{f}}_{t})=q({\vec{\mathbf{f}}},{\vec{\mathbf{u}}})q({\vec{\mathbf{x}}}),$ (42)
the factorization is known as a mean-field approximation in the GPSSM literature. Conversely, if the variational distribution, $q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}},{\vec{\mathbf{u}}})$, explicitly builds the dependence between the latent states and the transition function values, as shown in Eq. (8), it is a non-mean-field approximation.
Supplement D Variational Bayes
D-A Evidence Lower Bound (ELBO)
In Bayesian statistics, the model marginal likelihood $p({\vec{\mathbf{y}}}|{{\bm{\theta}}})$ is a fundamental quantity for model selection and comparison [38]. By maximizing the logarithm of $p({\vec{\mathbf{y}}}|{{\bm{\theta}}})$ w.r.t. the model parameters ${{\bm{\theta}}}$, the goodness of data fitting and the model complexity are automatically balanced, in accordance with Occam’s razor principle [37]. However, $p({\vec{\mathbf{y}}}|{{\bm{\theta}}})$ is obtained by integrating out all the latent variables $\{{\vec{\mathbf{x}}},{\vec{\mathbf{f}}}\}$ in the joint distribution, see Eq. (15), which is analytically intractable. Thus, the posterior distribution of the latent variables, $p({\vec{\mathbf{x}}},{\vec{\mathbf{f}}}|{\vec{\mathbf{y}}})=\frac{p({\vec{\mathbf{y}}},{\vec{\mathbf{x}}},{\vec{\mathbf{f}}})}{p({\vec{\mathbf{y}}}|{{\bm{\theta}}})},$ cannot be expressed in a closed-form expression, either. This intractability issue has been addressed in variational Bayesian methods by adopting a variational distribution [10], $q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})$, to approximate the intractable $p({\vec{\mathbf{x}}},{\vec{\mathbf{f}}}|{\vec{\mathbf{y}}})$. With the newly introduced variational distribution $q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})$, we have
$\!\!\!\log p({\vec{\mathbf{y}}}|{{\bm{\theta}}})=\log\frac{p({\vec{\mathbf{y}}},{\vec{\mathbf{x}}},{\vec{\mathbf{f}}})}{p({\vec{\mathbf{x}}},{\vec{\mathbf{f}}}|{\vec{\mathbf{y}}})}$ (43)
$\quad=\int\int q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})\log\frac{p({\vec{\mathbf{y}}},{\vec{\mathbf{x}}},{\vec{\mathbf{f}}})q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})}{p({\vec{\mathbf{x}}},{\vec{\mathbf{f}}}|{\vec{\mathbf{y}}})q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})}\mathrm{d}{\vec{\mathbf{x}}}\mathrm{d}{\vec{\mathbf{f}}}$
$\quad=\underbrace{\mathbb{E}_{q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})}\!\left[\log\frac{p({\vec{\mathbf{y}}},{\vec{\mathbf{x}}},{\vec{\mathbf{f}}})}{q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})}\right]}_{\text{ELBO: }\mathcal{L}}+\underbrace{\mathbb{E}_{q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})}\!\left[\log\frac{q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})}{p({\vec{\mathbf{x}}},{\vec{\mathbf{f}}}|{\vec{\mathbf{y}}})}\right]}_{\text{KL divergence}}$
Supplement E More Experiment Results
E-A OEnVI: Online Learning Results on Linear Gaussian SSMs
  • The definition of RMSE:
    $\text{RMSE}=\sqrt{\frac{1}{T}\sum_{t=1}^{T}\sum_{d=1}^{d_{x}}(\hat{{\mathbf{x}}}_{t}^{(d)}-{\mathbf{x}}_{t}^{(d)})^{2}}$ (44)
    where $\hat{{\mathbf{x}}}_{t}$ represents the estimation of ${\mathbf{x}}_{t}$.
  • The baseline in Table III is the RMSE between the noisy observations ${\mathbf{y}}_{1:T}$ and the latent states ${\mathbf{x}}_{1:T}$
TABLE III: State inference performance (RMSE) of EnVI and OEnVI
Time Slot 0 – 120 120 – 240 240 – 360 360–480 480–600 600–720 720–840 840–960 900-1000 0–1000
KF 0.5252 0.5149 0.5228 0.5658 0.5246 0.4907 0.4983 0.5202 0.4846 0.5199
EnVI 0.6841 / / / / / / / / 0.7182
OEnVI 0.7784 0.7130 0.6512 0.6487 0.6786 0.6515 0.5958 0.6713 0.6418 0.6739
Baseline 0.9872 0.9811 0.9510 1.0077 1.0215 0.9967 1.0077 1.0314 1.0222 0.9974
(a) Online learning result from $t=0$ to $t=120$
(b) Online learning result from $t=120$ to $t=240$
(c) Online learning result from $t=240$ to $t=360$
(d) Online learning result from $t=360$ to $t=480$
(e) Online learning result from $t=480$ to $t=600$
Figure 6: OEnVI on learning and inference in linear Gaussian SSMs
(a) Online learning result from $t=600$ to $t=720$
(b) Online learning result from $t=720$ to $t=840$
(c) Online learning result from $t=840$ to $t=960$
(d) Online learning result from $t=900$ to $t=1000$
(e) Online learning result from $t=0$ to $t=1000$
Figure 7: OEnVI on learning and inference in linear Gaussian SSMs
E-B Learning System Dynamics on Kink Function
  • The kink function is depicted in Fig. 8
  • The MSE and Log-likelihood in Table I are evaluated as follows:
    $\text{MSE}=\frac{1}{T}\sum_{t=1}^{T}\sum_{d=1}^{d_{x}}(\hat{{\mathbf{f}}}_{t}^{(d)}-{\mathbf{f}}_{t}^{(d)})^{2}$ (45)
[…]
Figure 8: The ”kink” dynamical function is used to generate $50$ latent states and corresponding observations, with $\sigma_{\mathrm{Q}}^{2}=0.01$ and $\sigma_{\mathrm{R}}^{2}=0.1$.
E-C Time Series Data Forecasting using EnVI
In addition to the overall prediction performance outlined in Table II, we provide a specific prediction of EnVI below.
(a) Actuator
(b) Ball Beam
(c) Drive
(d) Dryer
(e) Gas Furnace
Figure 9: Learning and prediction results of EnVI. The initial half of the sequence is generated by passing the filtered ${\mathbf{x}}_{t}$ through the emission model, while the subsequent half of the sequence represents the prediction outcome. This prediction is derived from the filtered ${\mathbf{x}}_{t}$ obtained from the final step of the training sequence, serving as the initial state.
E-D Online Learning and Inference with OEnVI
We report the learning results of OEnVI on the kink function dataset. As illustrated in Fig. 10, after sequentially training with $600$ data points, OEnVI demonstrates comparable performance to EnVI, which undergoes offline training consisting of over $400$ iterations with a full batch of data with length $T\!=\!600$.
Figure 10: Kink transition function learning using Online EnVI